# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ III.1.2 $T\to DP$ decay amplitudes
$T^{\prime\prime+}_{c}\to D^{0}\eta$ $\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{12}(F_{DP}+4AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}s_{\phi^{\prime}}-2c_{\phi^{\prime}}}{12}(F_{DP}+4AC_{P})$
$+\frac{2c_{\phi^{\prime}}-\sqrt{2}s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$ $+\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
@@
$T^{0}_{c}\to D^{0}\eta$ $\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{12}(F_{DP}+4AC_{P})$ $T^{0}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}s_{\phi^{\prime}}-2c_{\phi^{\prime}}}{12}(F_{DP}+4AC_{P})$
$+\frac{2c_{\phi^{\prime}}-\sqrt{2}s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$ $+\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $-\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{6}(F_{DP}+4AC_{P})$ $T^{+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{2c_{\phi^{\prime}}-\sqrt{2}s_{\phi^{\prime}}}{6}(F_{DP}+4AC_{P})$
$+\frac{2c_{\phi^{\prime}}-\sqrt{2}s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$ $+\frac{\sqrt{2}c_{\phi^{\prime}}+2s_{\phi^{\prime}}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
With the above effective Hamiltonians, we can decompose $T\to DP$ decay amplitudes in terms of these topological amplitudes. The $T\to DP$ decay amplitudes for $T=T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ in scenario I and $T=T(cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\})$ in scenario II are shown in Table 3 and 4, respectively. Some of the modes may be kinematically prohibited, but the information on these amplitudes can still be useful, as the tetraquark can be a virtual particle in some other processes.
@@
In fact, for flavor exotic states in scenario I, most of their decay amplitudes are identical up to a sign. As one can infer from Table 3, the squares of the amplitudes of the following six modes, namely $T^{++}_{c\bar{s}}\to D^{+}K^{+},D_{s}^{+}\pi^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$, $T^{0}_{c\bar{s}}\to D^{0}K^{0},D_{s}^{+}\pi^{-}$ and $T^{0}_{cs}\to D^{0}\overline{K}^{0},D^{+}K^{-}$ decays, should be identical, while those in $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ and $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ decays are half the above. The flavor exotic nature of most of these states is self-evident from the table. Note that in $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$ decay the exotic quantum number of $T^{+}_{c\bar{s}}$ can also be easily identified.
@@
The three other states in ${\bf 6}$, namely, $T^{+}_{c\bar{s}}$, $T^{+}_{c}$ and $T^{0}_{c}$, have $T\to DP$ decay rates that are proportional to the previous ones with simple factors. The situation is different for the three states in $\bar{\bf 3}$, namely $T^{\prime\prime+}_{c}$, $T^{\prime\prime 0}_{c}$ and $T^{\prime\prime+}_{cs}$, as their amplitudes contain $F_{DP}$, $AC_{P}$ and $AC_{D}$ as well. Nevertheless, the $T\to DP$ decay amplitudes of these states for $P=\pi$, $K$ and $\overline{K}$, are proportional to $F_{DP}+2AC_{P}$, giving $\Gamma(T\to D\pi^{+})=\Gamma(T\to D\overline{K})=\Gamma(T\to DK)=2\Gamma(T\to D\pi^{0})$ for these states. As noted previously we have $\bar{\bf 3}\otimes{\bf 8}=\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3}$, see Eq. (134). The $\bar{\bf 3}$ in the $D\Pi$ final states can match the $\bar{\bf 3}$ of $T$ in scenario I, but the addition of $\eta_{1}$ in $P$ provides another $\bar{\bf 3}$ in $DP$ and another combination, namely $F_{DP}-3AC_{D}-AC_{P}$, is needed for the $T\to D\eta,D\eta^{\prime}$ decay amplitudes.
@@
Similarly, for scenario II, as one can infer from Table 4, the following modes, namely, $T^{++}_{c\bar{s}}\to D^{+}K^{+},D_{s}^{+}\pi^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$, $T^{0}_{c\bar{s}}\to D^{0}K^{0},D_{s}^{+}\pi^{-}$ and $T^{0}_{cs}\to D^{0}\overline{K}^{0},D^{+}K^{-}$ decays, should have identical amplitudes squared. The amplitudes squared of $T^{+}_{cs}\to D^{+}\overline{K}^{0}$, $T^{-}_{cs}\to D^{0}K^{-}$, $T^{++}_{c\bar{s}\bar{s}}\to D_{s}^{+}K^{+}$, $T^{+}_{c\bar{s}\bar{s}}\to D_{s}^{+}K^{0}$, $T^{++}_{c}\to D^{+}\pi^{+}$ and $T^{-}_{c}\to D^{0}\pi^{-}$ decays are doubled compared to the former modes. In addition, for the other two flavor exotic states, we have $T^{+}_{c}\to D^{0}\pi^{+},D^{+}\pi^{0}$, $T^{0}_{c}\to D^{0}\pi^{0}$ and $D^{+}\pi^{-}$ decays with amplitudes squared porpotianal to the aboves. The flavor exotic nature of most of these states is self-evident from the table. For $T^{+}_{c\bar{s}}$, the $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$ decay reveals the exotic nature of $T^{+}_{c\bar{s}}$, but for $T^{+}_{c}$ and $T^{0}_{c}$, elebrated works are required, as one needs to verify that they do not decay to $D\eta$ and $D\eta^{\prime}$.
@@
Furthermore, four other states in $\overline{\bf 15}$, namely, $T^{+}_{c\bar{s}}$, $T^{\prime+}_{c}$, $T^{\prime 0}_{c}$ and $T^{\prime+}_{c\bar{s}}$, have $T\to DP$ decay amplitudes squared proportional to the previous ones with some simple factors, while the situation is slightly different for the three states in $\bar{\bf 3}^{\prime}$, namely $T^{\prime\prime+}_{c}$, $T^{\prime\prime 0}_{c}$ and $T^{\prime\prime+}_{cs}$, as their amplitudes contain $F_{DP}$, $AC_{P}$ and $AC_{D}$ as well. Nevertheless, as shown in Table 4, the $T\to DP$ decay amplitudes of these states for $P=\pi$, $K$ and $\overline{K}$, are proportional to the combination of $F_{DP}+4AC_{P}$, giving $\Gamma(T\to D\pi^{+})=\Gamma(T\to D\overline{K})=\Gamma(T\to DK)=2\Gamma(T\to D\pi^{0})$ for these states. The $\bar{\bf 3}$ in the $D\Pi$ final states can match the $\bar{\bf 3}^{\prime}$ of $T$ in scenario II, but the addition of $\eta_{1}$ in $P$ provide another $\bar{\bf 3}$ in $DP$, where another combination, namely $3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})$, is needed for the $T\to D\eta,D\eta^{\prime}$ decay amplitudes.

### v3

@@ 6 $T^{0}_{c}\to D^{0}\pi^{0}$ $\frac{1}{2}F_{DP}$ $T^{0}_{c}\to D^{+}\pi^{-}$ $\frac{F_{DP}}{\sqrt{2}}$
$T^{\prime\prime+}_{c}\to D^{0}\eta$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{12}(F_{DP}+4AC_{P})$ $T^{\prime\prime+}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}s_{\theta}-2c_{\theta}}{12}(F_{DP}+4AC_{P})$
$\frac{2c_{\theta}-\sqrt{2}s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})] +$$\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
@@
$T^{0}_{c}\to D^{0}\eta$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{12}(F_{DP}+4AC_{P})$ $T^{0}_{c}\to D^{+}\eta^{\prime}$ $\frac{\sqrt{2}s_{\theta}-2c_{\theta}}{12}(F_{DP}+4AC_{P})$
$\frac{2c_{\theta}-\sqrt{2}s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})] +$$\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\eta$ $\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{6}(F_{DP}+4AC_{P})$ $T^{+}_{c\bar{s}}\to D_{s}^{+}\eta^{\prime}$ $\frac{2c_{\theta}-\sqrt{2}s_{\theta}}{6}(F_{DP}+4AC_{P})$
$\frac{2c_{\theta}-\sqrt{2}s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})] +$$\frac{\sqrt{2}c_{\theta}+2s_{\theta}}{3}[3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})]$
With the above effective Hamiltonians, we can decompose $T\to DP$ decay amplitudes in terms of these topological amplitudes. The $T\to DP$ decay amplitudes for $T=T(cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}])$ in scenario I and $T=T(cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\})$ in scenario II are shown in Table 3 and 4, respectively. Note that some of the modes may be kinematically prohibited, but the information on these amplitudes can still be useful, as the open-charmed tetraquark can be a virtual particle in some other processes.
@@
In fact, for flavor exotic states in scenario I, most of their decay amplitudes are identical up to a sign. As one can infer from Table 3, the squares of the amplitudes of the following six modes, namely $T^{++}_{c\bar{s}}\to D^{+}K^{+},D_{s}^{+}\pi^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$, $T^{0}_{c\bar{s}}\to D^{0}K^{0},D_{s}^{+}\pi^{-}$ and $T^{0}_{cs}\to D^{0}\overline{K}^{0},D^{+}K^{-}$ decays, should be identical, while those in $T^{+}_{c\bar{s}}\to D^{0}K^{+}$ and $T^{+}_{c\bar{s}}\to D^{+}K^{0}$ decays are half the above. The flavor exotic nature of most of these states is self-evident from their decay modes shown in the table.
The three other states in $\bf 6$}, namely, $T^{+}_{c\bar{s}}$, $T^{+}_{c}$ and $T^{0}_{c}$, have $T\to DP$ decay rates that are proportional to the previous ones with simple factors. Note that in $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$ decay the exotic quantum number of $T^{+}_{c\bar{s}}$ can also be easily identified, as $\pi^{0}$ is an iso-vector, $T^{+}_{c\bar{s}}$ cannot be a $c\bar{s}$ state.
@@
The situation is different for the three states in $\bar{\bf 3}$, namely $T^{\prime\prime+}_{c}$, $T^{\prime\prime 0}_{c}$ and $T^{\prime\prime+}_{cs}$, as their amplitudes contain $F_{DP}$, $AC_{P}$ and $AC_{D}$ as well. Nevertheless, the $T\to DP$ decay amplitudes of these states for $P=\pi$, $K$ and $\overline{K}$, are proportional to $F_{DP}+2AC_{P}$, giving $\Gamma(T\to D\pi^{+})=\Gamma(T\to D\overline{K})=\Gamma(T\to DK)=2\Gamma(T\to D\pi^{0})$ for these states. As noted previously we have $\bar{\bf 3}\otimes{\bf 8}=\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3}$, see Eq. (176). The $\bar{\bf 3}$ in the $D\Pi$ final states can match the $\bar{\bf 3}$ of $T$ in scenario I, but the addition of $\eta_{1}$ in $P$ provides another $\bar{\bf 3}^{\prime}$ in $DP$ and another combination, namely $F_{DP}-AC_{P}-3AC_{D}$, is needed for the $T\to D\eta,D\eta^{\prime}$ decay amplitudes.
Indeed, the above form on $T\to DP$ decay amplitudes still holds in the presence of rescattering, as strong interaction does not affect the flavor structure of the decay amplitudes. For illustration, we consider the effect of final state interaction in the $DP$ system with quasi-elastic strong rescattering Chua:2001br ; Chua:2005dt ; Chua:2007qw . Before rescattering the topological amplitudes in the $T\to DP$ decay amplitudes are given by Table 3, but with $F_{DP}$, $F_{DP}+2AC_{P}$ and $F_{DP}-AC_{P}-3AC_{D}$ replaced with $F^{0}_{DP}$, $F^{0}_{DP}+2AC^{0}_{P}$ and $F^{0}_{DP}-AC^{0}_{P}-3AC^{0}_{D}$ (superscript 0 are added to indicate amplitudes before rescattering), respectively. The question is whether the form of $T\to DP$ decay amplitudes, in terms of topological amplitudes, holds in the presence of the rescattering. As shown in Appendix B, after rescattering, the form of the decay ampltudes remain unchanged, but with topological amplitudes modified and redefined as follows,
$F_{DP}$ $=$ $e^{i\delta_{6}}F^{0}_{DP},$
$\left(\begin{array}[]{c}F_{DP}+2AC_{P}\\ F_{DP}-AC_{P}-3AC_{D}\end{array}\right)$ $=$ $\left(\begin{array}[]{cc}e^{i\delta_{\bar{3}}}\cos^{2}\tau+e^{i\delta_{\bar{3}^{\prime}}}\sin^{2}\tau&\frac{e^{i\delta_{\bar{3}}}-e^{i\delta_{\bar{3}^{\prime}}}}{\sqrt{2}}\sin\tau\cos\tau\\ \sqrt{2}(e^{i\delta_{\bar{3}}}-e^{i\delta_{\bar{3}^{\prime}}})\sin\tau\cos\tau&e^{i\delta_{\bar{3}^{\prime}}}\cos^{2}\tau+e^{i\delta_{\bar{3}}}\sin^{2}\tau\end{array}\right)$ (33)
$\times\left(\begin{array}[]{c}F^{0}_{DP}+2AC^{0}_{P}\\ F^{0}_{DP}-AC^{0}_{P}-3AC^{0}_{D}\end{array}\right),$
where $\delta_{6}$, $\delta_{\bar{3}}$ and $\delta_{\bar{3}^{\prime}}$ are strong phases and $\tau$ is the mixing angle between the $\bar{\bf 3}$ and $\bar{\bf 3}^{\prime}$ in the $DP$ system. For example, according to Table 3, before rescattering the $T^{++}_{c\bar{s}}\to D^{+}K^{+}$ and $D_{s}^{+}\pi^{+}$ amplitudes are $F^{0}_{DP}$ and $F^{0}_{DP}$, respectively. As shown in Appendix B, after quasi-elastic strong rescattering, they become $e^{i\delta_{6}}F^{0}_{DP}$ and $e^{i\delta_{6}}F^{0}_{DP}$, respectively, which agree to those shown in Table 3 by taking $F_{DP}=e^{i\delta_{6}}F^{0}_{DP}$. Therefore, the form of the amplitudes in terms of topological amplitudes is unchanged, but the topological amplitude itself is affected by rescattering. This agrees with the common understanding that topological amplitudes can include long-distance final state interaction effects, see, for example, Chau:tk ; Chau:1990ay ; Rosner:1999xd ; Cheng:2002ai .
@@
Similarly, for scenario II, as one can infer from Table 4, the following modes, namely, $T^{++}_{c\bar{s}}\to D^{+}K^{+},D_{s}^{+}\pi^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$, $T^{0}_{c\bar{s}}\to D^{0}K^{0},D_{s}^{+}\pi^{-}$ and $T^{0}_{cs}\to D^{0}\overline{K}^{0},D^{+}K^{-}$ decays, should have identical amplitudes squared. The amplitudes squared of $T^{+}_{cs}\to D^{+}\overline{K}^{0}$, $T^{-}_{cs}\to D^{0}K^{-}$, $T^{++}_{c\bar{s}\bar{s}}\to D_{s}^{+}K^{+}$, $T^{+}_{c\bar{s}\bar{s}}\to D_{s}^{+}K^{0}$, $T^{++}_{c}\to D^{+}\pi^{+}$ and $T^{-}_{c}\to D^{0}\pi^{-}$ decays are doubled compared to the former modes. In addition, for the other two flavor exotic states, we have $T^{+}_{c}\to D^{0}\pi^{+},D^{+}\pi^{0}$, $T^{0}_{c}\to D^{0}\pi^{0}$ and $D^{+}\pi^{-}$ decays with amplitudes squared porpotianal to the aboves. The flavor exotic nature of most of these states is self-evident from their decay modes shown in the table. For $T^{+}_{c\bar{s}}$, the $T^{+}_{c\bar{s}}\to D_{s}^{+}\pi^{0}$ decay reveals the exotic nature of $T^{+}_{c\bar{s}}$, but for $T^{+}_{c}$ and $T^{0}_{c}$, elebrated works are required, as one needs to verify that they do not decay to $D\eta$ and $D\eta^{\prime}$.
@@
Furthermore, four other states in $\overline{\bf 15}$, namely, $T^{+}_{c\bar{s}}$, $T^{\prime+}_{c}$, $T^{\prime 0}_{c}$ and $T^{\prime+}_{c\bar{s}}$, have $T\to DP$ decay amplitudes squared proportional to the previous ones with some simple factors, while the situation is slightly different for the three states in $\bar{\bf 3}^{\prime}$, namely $T^{\prime\prime+}_{c}$, $T^{\prime\prime 0}_{c}$ and $T^{\prime\prime+}_{cs}$, as their amplitudes contain $F_{DP}$, $AC_{P}$ and $AC_{D}$ as well. Nevertheless, as shown in Table 4, the $T\to DP$ decay amplitudes of these states for $P=\pi$, $K$ and $\overline{K}$, are proportional to the combination of $F_{DP}+4AC_{P}$, giving $\Gamma(T\to D\pi^{+})=\Gamma(T\to D\overline{K})=\Gamma(T\to DK)=2\Gamma(T\to D\pi^{0})$ for these states. The $\bar{\bf 3}$ in the $D\Pi$ final states can match the $\bar{\bf 3}^{\prime}$ of $T$ in scenario II, but the addition of $\eta_{1}$ in $P$ provide another $\bar{\bf 3}^{\prime}$ in $DP$, where another combination, namely $3AC_{D}+\sqrt{2}(F_{DP}+AC_{P})$, is needed for the $T\to D\eta,D\eta^{\prime}$ decay amplitudes.
Similarly, under quasi-elastic scattering Chua:2001br ; Chua:2005dt ; Chua:2007qw , the topological amplitudes are given by, see Appendix B,
$F_{DP}$ $=$ $e^{i\delta_{\overline{15}}}F^{0}_{DP},$
$\left(\begin{array}[]{c}F_{DP}+4AC_{P}\\ F_{DP}+AC_{P}+\frac{3}{\sqrt{2}}AC_{D}\end{array}\right)$ $=$ $\left(\begin{array}[]{cc}e^{i\delta_{\bar{3}}}\cos^{2}\tau+e^{i\delta_{\bar{3}^{\prime}}}\sin^{2}\tau&\sqrt{2}(e^{i\delta_{\bar{3}^{\prime}}}-e^{i\delta_{\bar{3}}})\sin\tau\cos\tau\\ \frac{e^{i\delta_{\bar{3}^{\prime}}}-e^{i\delta_{\bar{3}}}}{\sqrt{2}}\sin\tau\cos\tau&e^{i\delta_{\bar{3}^{\prime}}}\cos^{2}\tau+e^{i\delta_{\bar{3}}}\sin^{2}\tau\end{array}\right)$ (41)
$\times\left(\begin{array}[]{c}F^{0}_{DP}+4AC^{0}_{P}\\ F^{0}_{DP}+AC^{0}_{P}+\frac{3}{\sqrt{2}}AC^{0}_{D}\end{array}\right),$
where $\delta_{\overline{15}}$, $\delta_{\bar{3}}$ and $\delta_{\bar{3}^{\prime}}$ are strong phases and $\tau$ is the mixing angle between the $\bar{\bf 3}$ and $\bar{\bf 3}^{\prime}$ in the $DP$ system, and $F^{0}_{DP}$, $F^{0}_{DP}+4AC^{0}_{P}$ and $F^{0}_{DP}+AC^{0}_{P}+\frac{3}{\sqrt{2}}AC^{0}_{D}$ are the topological amplitudes before rescattering. The form of $T\to DP$ decay amplitudes in terms of the topological amplitudes does not change under rescattering. For example, according to Table 4, before rescattering the $T^{++}_{c\bar{s}}\to D^{+}K^{+}$ and $D_{s}^{+}\pi^{+}$ amplitudes are $F^{0}_{DP}$ and $F^{0}_{DP}$, respectively. As shown in Appendix B, after quasi-elastic strong rescattering, they become $e^{i\delta_{\overline{15}}}F^{0}_{DP}$ and $e^{i\delta_{\overline{15}}}F^{0}_{DP}$, respectively, which agree to those shown in Table 4 by taking $F_{DP}=e^{i\delta_{\overline{15}}}F^{0}_{DP}$. Therefore, the form of the amplitudes in terms of topological amplitudes is unchanged, but the topological amplitude itself is affected by rescattering.